A contractor mixes concrete from bags of pre-mix for small jobs. How many bags with 10% cement should he mix with 8 bags of 12.5% cement to produce a mix containing 11% cement.
Found 2 solutions by nyc_function, modi.kk91:
Answer by nyc_function(2741)

Let b = bags of cement
0.10b + 8(0.125) = 0.11(b + 8)
Can you take it from here?
I gave you the equation that you'll need to find the answer.

Answer by modi.kk91(20)

solved by allegation method
10% 12.5%


11%

12.5-11=1.5 and 11-10=1

ratio=15:10=3:2

3:2::x:8


x=12